If something was $0.40 and its increased to $0.90 what is the precent increase?
Answer by stanbon(75887)

If something was $0.40 and its increased to $0.90 what is the precent increase?
------
Quantity Increase: 90-40 = 50
----
Percent increase: 50/40 = 5/4 = 1.25 = 125%
------------------------------------------------
Cheers,
Stan H.
======================